OpenStudy (debbieg):

New users need a "How to best utilize OS" tip sheet, maybe incorporated into the CoC, or as a separate item.

OpenStudy (debbieg):

I'm talking about things like: - Post each question separately, don't just keep adding more problems in the same thread once the first one has been answered. - The people answering your questions are volunteering their time to help you. Be sure to acknowledge them, say thank you, and award a medal for the answer that you found most helpful. {I've worked to type up elaborate answers, and not gotten so much as a thank-you or acknowledgement that it as even seen.... annoying!} - Don't private-message questions. If you would like a particular person who has helped you in the past to look at your question, just tag them in the thread. I'm sure there are more..... a bit beyond the scope of the current CoC, but I think would be helpful to many new (and some older) users.

HanAkoSolo (jamierox4ev3r):

what we need is a way for the users to actually read the Code of Conduct also !!! and also it looks like you are basically trying to add to the Code of Conduct. Why don't you contact a moderator on this site?

OpenStudy (debbieg):

and I have noticed that some of the ambies are mentioning medals and directing new users to the CoC - I think that's great!! I just think there are other things that could be "available" somewhere that they could refer to.

HanAkoSolo (jamierox4ev3r):

this is an imperfect system.. but we are doing everthing we can :)

thomaster (thomaster):

Great idea :) A FAQ page would be nice too (frequently asked questions) which covers many of the daily questions like: how to become ambi/mod?, or does openstudy have an app?

OpenStudy (debbieg):

Yes, I'm not trying to be critical... just making a suggestion. :) I don't really see these as things that fall under "code of conduct" issues, more like "how to get the most out of this site" kinds of things. YES, @thomaster , something along the lines of a FAQ!

terenzreignz (terenzreignz):

Compile some :) LOL PM questions usually imply somebody seeking only the answer. Or an under-the-table deal. Toss up :P

OpenStudy (e.mccormick):

And an official \(\LaTeX\) tutorial. \(\ddot\smile\)

terenzreignz (terenzreignz):

You got something against smileys? D: :)

thomaster (thomaster):

\(\Huge ☺\)

terenzreignz (terenzreignz):

Where IS that Latex Practicing :) study group anyway... (and yes, that smiley :) is part of the name...)

OpenStudy (missmob):

pies

OpenStudy (e.mccormick):

@terenzreignz Nothing at all against them. In fact, I am leading the \(\LaTeX\) charge here to expand from just =) to simple things like: \[ \,\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \hat\odot \; \hat\odot \\ \Downarrow\\ \Large \color{#CD5555}{\smile} \end{matrix} \right)}}\normalsize \\ \begin{array}{lcr} /&\bowtie&\backslash\\ |&\scriptsize^\circ&|\\ |&\scriptsize^\circ&|\\ \end{array} \] That way people have options. Do they want to wink, ;-) , or do they want to inspire! \(\Huge \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\bullet \widehat\bullet} \\ \color{red}\smile\end{matrix}\right)}}\)

terenzreignz (terenzreignz):

^show off -.- LOL Just kidding :D

OpenStudy (e.mccormick):

Well the bottom one is DebbieG's creation, so I felt it was only proper to show it in her post here!

OpenStudy (debbieg):

Haha... I need to copyright it!\[\Huge \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\ \color{red}\smile\end{matrix}\right)}}_{\Large©debbieg}\]

OpenStudy (e.mccormick):

\[\Huge \color{violet}{\overbrace{ \left( \begin{matrix}\color{blue}{\widehat\otimes \widehat\otimes} \\ \color{red}\frown\end{matrix}\right)}}\]